# REPS: Recycled Entropy Packet Spraying for Adaptive Load Balancing and Failure Mitigation

### v5

@@ Acknowledgements.
We thank the Swiss National Supercomputing Center (CSCS) for providing the computational resources used in this work. The authors used ChatGPT to only assist with editing and quality control throughout this manuscript (ideas and content remain original).
@@ Appendix A Freezing Mode in REPS
Regardless of the employed strategy, REPS maintains high performance even if it inadvertently enters freezing mode without an actual network failure. This is because entering freezing mode effectively reduces the EVS size of REPS. As demonstrated in Section 4.5.2, REPS remains effective with as few as 32 EVs. For instance, we tested REPS with the 16 MiB tornado workload, running it first under normal conditions and then with forced freezing mode activated after $150\;\mu\text{s}$. The results showed only a 1% increase in completion time for REPS with forced freezing mode, with the 99th percentile RTT remaining very close, as illustrated in Figure 15.
@@
Figure 15. Studying REPS while forcing freezing mode without failures.
Appendix B Entropy Values Set Analysis
Practical implementations make use of hashing functions to map EVs onto output ports. We show that while using a small EVS can lead to high inherent load balancing issues, using $2^{16}$ EVs is very close to uniformly random.
We first show theoretically that the EVS size is very important for OPS and later that REPS is much more flexible and can work well even with small EVS.
We use a balls-into-bins model (Raab and Steger, 1998), where $m$ balls are thrown uniformly and idependently at random into $n$ bins. The goal is to determine the largest number of balls in any bin, referred to as the maximum load $l(m,n)$. We define the load imbalance $\lambda_{m,n}$ as $\frac{l(m,n)}{m/n}-1$, representing the extent to which the most heavily loaded bin exceeds the average.
We confirm this theoretical analysis with simulations of the load imbalance. Figures 16(a) and 16(b) show the distribution of the load imbalance for $1$ and $32$ unique flows, respectively. We note that each flow is from a different sender and will hence have different header fields that will be used in the hashing function regardless of the EV value. We note that for each case, we throw for each active flow a number of balls equal to the EVS size, with each ball being a unique EV. We can see that for 32 flows, choosing less than $2^{8}$ EVs can lead to more than $10\%$ load-imbalance, whereas $2^{16}$ EVs guarantee less than a $1\%$ load imbalance.
((a)) 1 flow active
((b)) 32 flows active
Figure 16. The expected load imbalance at a switch with 32 uplinks.
Appendix C REPS Single Switch Convergence Proof
@@
where we show the tail bound in Lemma C.2.
@@
Next, we bound the expected drift when there is more than one ball per color. We have two ball of the same color only if a ball does not get removed before the next ball of the same color is thrown. This happens if more than $b$ balls arrive in the same bin in the same time step as the ball (this uses the FiFO-property of the queues). By a tail bound on the binomial distribution, this happens with probability at most $\frac{e^{5/4}}{n^{3}}$, by Lemma C.1 using $b\geq\frac{12}{5}\ln n$. By a union bound, the probability that any color in a batch has more than one ball is at most $\frac{e^{5/4}}{n^{2}}$. The drift is bounded by $n$, so the contribution to the expected drift by this case is at most $\frac{e^{5/4}}{n}$.
@@
$E\left[B\ |\ B\geq x\right]$ $\leq x+\frac{(n-x)\frac{1}{n}}{x-\frac{k}{n}+\frac{1}{n}}\leq x+\frac{1}{x-1}\enspace.$
We bound $P[B\geq x]$ similarly as in Lemma C.1:
@@
Appendix D Additional Results
@@
D.1. ACK Coalescing Theoretical Modeling
@@ D.3. Incremental Failures
Appendix E Additional Data

### v6

@@ Acknowledgements.
We thank our shepherd, Daniel Amir, and the anonymous reviewers of EuroSys ’26 for their valuable feedback. This work is supported by the European Union’s Horizon Europe under grant 101175702 (NET4EXA), the Sapienza University Grants ADAGIO and D2QNeT (Bando per la ricerca di Ateneo 2023 and 2024), the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement PSAP, No. 101002047), and a CAF America grant. We thank the Swiss National Supercomputing Center (CSCS) for providing the computational resources used in this work. We also thank Marcin Copik, Ales Kubicek, and Nadeen Gebara for providing their valuable feedback. We used ChatGPT solely for text editing and quality checks.
@@ Appendix A Freezing Mode in REPS
Regardless of the employed strategy, REPS maintains high performance even if it inadvertently enters freezing mode without an actual network failure. This is because entering freezing mode effectively reduces the EVS size of REPS. As demonstrated in Section 4.5.2, REPS remains effective with as few as 32 EVs. For instance, we tested REPS with the 16 MiB permutation workload, running it first under normal conditions and then with forced freezing mode activated after $50\;\mu\text{s}$. The results show comparable results for REPS with or without forcing the freezing mode, as illustrated in Figure 19 with only minor instability introduced by activating freezing mode. However, compared to OPS, REPS with the force freezing mode is still significantly more stable and has a finishing time comparable to the standard REPS, both completing the workload slightly faster than OPS.
@@
Figure 19. Studying REPS while forcing freezing mode after a fixed period of time.
Appendix B REPS Single Switch Convergence Proof
@@
where we show the tail bound in Lemma B.2.
@@
Next, we bound the expected drift when there is more than one ball per color. We have two balls of the same color only if a ball does not get removed before the next ball of the same color is thrown. This happens if more than $b$ balls arrive in the same bin in the same time step as the ball (this uses the FiFO-property of the queues). By a tail bound on the binomial distribution, this happens with probability at most $\frac{e^{5/4}}{n^{3}}$, by Lemma B.1 using $b\geq\frac{12}{5}\ln n$. By a union bound, the probability that any color in a batch has more than one ball is at most $\frac{e^{5/4}}{n^{2}}$. The drift is bounded by $n$, so the contribution to the expected drift by this case is at most $\frac{e^{5/4}}{n}$.
@@
$E\left[B\ |\ B\geq x\right]$ $\leq x+\frac{(n-x)\frac{1}{n}}{x-\frac{k}{n}+\frac{1}{n}}\leq x+\frac{1}{x-1}\kern 5.0pt.$
We bound $P[B\geq x]$ similarly as in Lemma B.1:
@@
Appendix C Additional Results
@@
C.1. ACK Coalescing Theoretical Modeling
@@ C.3. Incremental Failures
C.4. Impact of Freezing Mode
To demonstrate the impact of having freezing mode in REPS, we evaluate three different scenarios with and without freezing mode. In Figure 23, we show the effect of having freezing mode available versus disabling it. As expected, in scenarios without failures, REPS performs the same with or without freezing mode. However, when adding failures in the form of 1% cables failure, we can see freezing mode helping, with almost a 25% gain over not having it. However, we highlight that REPS remains competitive even when running without freezing mode. This could be used to simplify deployments if needed.
Figure 23. Impact on runtime of disabling freezing mode in REPS.
Appendix D Additional Data
@@
Appendix E Artifact Appendix
E.1. Abstract
This appendix provides concise instructions for installing and executing the REPS artifact. Each figure is generated by a dedicated Python script to ensure reproducibility. We also include Bash scripts to run the full experiment suite or any user-selected subset. Note that the AD was prepared earlier than the camera-ready paper, so some figure numbers may mismatch the camera-ready version.
E.2. Description & Requirements
E.2.1. How to access
We provide a link to a GitHub repository and a link to a Zenodo DOI:
In the content of the repository we also offer instructions inside the README.
E.2.2. Hardware dependencies
No specialized hardware is required. All experiments were performed on a commodity x86 CPU. Any similar processor capable of running the simulator should suffice to reproduce the results. For our runs we used 40GB of RAM, but we expect the simulations to run also with 32GB of RAM. 16GB of RAM might also work but might require lowering the number of parallel simulations.
E.2.3. Software dependencies
The artifact requires C++17 and Python3.8 as the minimum dependencies for building and running both the htsim simulator and the relative Python scripts. All experiments and computational artifacts in this paper were run locally on Ubuntu 22.04 LTS using WSL 2.
To process and visualize the raw data, we rely on a set of standard Python packages, i.e., seaborn, scipy, numpy, and pandas. We do provide a requirements.txt that users can use to install all needed Python dependencies.
E.2.4. Benchmarks
All the benchmarks are self contained and do not depend on external data.
E.3. Set-up
After obtaining the source code, either via git clone or by downloading and extracting the Zenodo archive, install the required packages by running:
  • ./reps_pkg_install.sh
If you encounter errors or want to avoid conflicts with local packages, run these commands inside a clean Python environment created with venv. For example:
  • python3 -m venv .venv
  • source .venv/bin/activate
Then, from the ./htsim/sim directory, install htsim with:
  • make clean && cd datacenter/ && make clean &&
    cd .. && make -j 8 && cd datacenter/ && make -j 8 &&
    cd ..
E.4. Evaluation workflow
E.4.1. Major Claims
Our main claim of the paper is that:
  • (C1): REPS consistently matches or exceeds the performance of modern load-balancing schemes. On symmetric networks, it always slightly outperforms OPS; under failures or asymmetric conditions, it delivers clear gains over OPS and other adaptive approaches.
E.4.2. Experiments
Please note that minor discrepancies from the submitted version may occur due to hardware or software differences and inherent randomness, despite our efforts to fix all random seeds. Additionally, some plots were manually annotated and enhanced during post-processing. However, for most figures we expect an exact match. With that in mind, running and analyzing the code is straightforward and explained in the next paragraph.
[Execution] We provide in the /artifact_scripts folder, a number of Python scripts that will run the code needed for their respective figures. For example, to generate Figure 1, one must run the fig_1_symmetric_micro.py Python script.
To facilitate the process, we provide three Bash scripts that can run most or all experiments needed to reproduce the results of the paper. In particular:
  • reps_quick.sh: runs quickly (less than 2 hours) and generates a small subset of Figures.
  • reps_medium.sh: runs in about 6 hours and generates almost all Figures of the paper.
  • reps_full.sh: runs all experiments in the paper and may take up to a day.
Exact running times depend on hardware performance.
[Results] Once experiments are run, results are automatically generated in the artifact_results/ directory. Within this directory, each experiment has its own subfolder containing the plots and, when applicable, the raw_data/.
For instance, to visualize the first Figure, one must navigate to artifact_results/fig_1_symmetric_micro/plots and open its contents. In most cases we provide the plots in both .pdf and .png format.